Today we worked on the exploration of the angular DOFs working points with the DRMI configuration (arms locked at -3kHz of CARM offset). As set yesterday, the CHRoCC setpoint was 0.43 V.
Preliminarly, we did a few checks:
- from 6:50 to 6:56 UTC we made a scan of the arm cavities in order to check the mismatch of the arms: the results are similar toSaturday's measurements, but there is a slight improvement on both arms: NArm is at ~2.8% while WArm is at 1.35%;
- around 7:20 UTC, we did a minor tuning of the DRMI demodulation phases (both moved by 0.1 rad);
- around 7:57 UTC, we found out that the occasional bursts of noise that lately have been seen in the DRMI configuration were originated in the CARM loop, which had too much gain; we lowered it by 40%, causing the noise to disappear and all the correction signals of the DRMI to be reduced;
- during this first lock of the DRMI (7:52 UTC for reference) the B4 phase camera imagess were pretty circular and symmetric as expected, while the B1p camera images were quite bad; in this working point, as reference, the main figures of merit were: B4_DC = 2.56, B4_12MHz_mag= 0.135; USB and LSB 56 MHz power respectively 12 and 8.
Then we moved to the topic of the shift, here are the measurements we performed; occasionally, we relocked from scratch in order to recover the pre-alignment of the arms:
- we started with a wide scan of the PR transversal position, first on the Y direction (starting at 8:06 UTC for the positive direction, 8:19 UTC for the negative one); at the very last step of the negative scan the DRMI unlocked;
- at 8:32 UTC we started the scan on PR_X (positive direction), while at 8:48 UTC the scan was performed in the negative direction; also in this case, at the very end of the latter the DRMI unlocked;
- at 09:01:30 UTC we started the scan of the PR angular working point (on TX; TY started at 9:14 UTC): given that we have a full bandwidth loop for this DOF, we scanned the loop's setpoints, looking at the usual 2D plots; here we focused on the maximization of the B4_12MHz_mag signal, and we used the setpoints we found here onwards; at this point (9.23 UTC) the superposition of 56MHz LSB and USB powers were respectively 8 and 14;
- at 10:18:30 UTC, after a full pre-alignment we relocked the DRMI locked; the superposition of 56MHz LSB and USB powers were respectively 8 and 13;
- at 10:30 UTC we opened the BS drift control and started the scan on the TX DOF, positive direction, with the local controls;
- we relocked the DRMI from scratch at 12:14 UTC; here the 56MHz LSB and USB powers were respectively 6 and 11;
- we started the scan on BS_TX negative direction at 12:19 UTC;
- at 13:07 UTC we relocked the arms in order to realign, then we relocked the DRMI; the 56MHz LSB and USB powers were respectively 6 and 12;
- we started the scan of BS_TY (positive) at 13:11 UTC, and BS_TY (negative) at 13:42 UTC; in both cases, after reaching a misaglinment of (-) 0.9 urad we witnessed the start of an oscillation on several signals, so we did not move any further;
- at 14:03:25 UTC we unlocked in order to realign the arms;
- finally, we scanned the SR DOFs using the local controls: at 14:23 UTC we started the scan on TX positive, then at 14:49 UTC TX negative; TY positive and negative scans started respectively at 14:59:26 and 15:18 UTC.
- 1) mismatch measurement;
- 2) to 5): scan of PR_F7_Y, phase camera images before the scan, after positive scan and negative scan respectively;
- 6) to 8): scan of PR_F0_X, phase camera images after positive scan and negative scan respectively;
- 9) to 11): scan of PR angular DOFs, and 2D plots;
- 12) to 15); scan of BS local controls, TX positive and negative, TY positive and negative respectively;
- 16) and 17): scan of SR local controls, TX and TY respectively.
Data will be analyzed offline, here just a preliminary comment: there doesn't seem to be a strong evidence about a preferred alignment position different from the standard, as basically all the figures of merit seem to degrade (sidebands and DC powers, dark fringe, PC overlap signals); more attention should be devoted to the study of SR, as the results for its DOFs are less clear, and the interaction with the BS drift control (which moved the mirror during the SR scans) should be studied more.